OpenStudy (anonymous):

David is rowing a boat upstream. The river is flowing at a speed of 2 miles per hour. David starts rowing at a speed of 6 miles per hour, but as he gets tired his speed decreases (at a rate of 1 mile per hour, every hour). Which equation represents the speed of the boat for x hours spent rowing?

OpenStudy (anonymous):

@Kemist

OpenStudy (anonymous):

is C. y = 4 − x an option ???

OpenStudy (anonymous):

y=4-x ? one of your options?

OpenStudy (anonymous):

> y = -x + C When x = 0, y = 6 - 2 => 4 so, 4 = C i.e. y = -x + 4 => 4 - x